### v2

# G-Transformer: Counterfactual Outcome Prediction under Dynamic and Time-varying Treatment Regimes

Abstract
In the context of medical decision making, counterfactual prediction enables clinicians to predict treatment outcomes of interest under alternative courses of therapeutic actions given observed patient history. Prior machine learning approaches for counterfactual predictions under time-varying treatments focus on static time-varying treatment regimes where treatments do not depend on previous covariate history. In this work, we present G-Transformer, a Transformer-based framework supporting g-computation for counterfactual prediction under dynamic and time-varying treatment strategies. G-Transfomer captures complex, long-range dependencies in time-varying covariates using a Transformer architecture. G-Transformer estimates the conditional distribution of relevant covariates given covariate and treatment history at each time point using an encoder architecture, then produces Monte Carlo estimates of counterfactual outcomes by simulating forward patient trajectories under treatment strategies of interest. We evaluate G-Transformer extensively using two simulated longitudinal datasets from mechanistic models, and a real-world sepsis ICU dataset from MIMIC-IV. G-Transformer outperforms both classical and state-of-the-art counterfactual prediction models in these settings. To the best of our knowledge, this is the first Transformer-based architecture for counterfactual outcome prediction under dynamic and time-varying treatment strategies. Code will be released upon publication of the paper.
1 Introduction
Clinicians often have to choose among multiple treatment options for their patients but do not have the ability to test every strategy before making a decision. Counterfactual prediction in medical decision-making involves the estimation of potential future trajectories of covariates of interest under alternative courses of action given observed history.
Treatment strategies of interest are typically time-varying, indicating decisions span multiple time points, and dynamic, implying that each treatment decision at a given time point is influenced by the preceding history up to that point. In selecting among competing dynamic treatment strategies, it is useful to obtain counterfactual predictions regarding a patient’s probability of experiencing adverse outcomes under each alternative strategy, based on their observed covariate history. Recent works presented deep learning approaches to estimate time-varying treatment effects (Lim et al., 2018; Bica et al., 2020a, b; Melnychuk et al., 2022) . However, most previous approaches focus on estimating counterfactual outcomes under static time-varying treatment strategies where treatments are not dependent on past covariate history.
In this work, we present G-Transformer, a transformer-based (Vaswani et al., 2017) framework for counterfactual prediction under dynamic and time-varying treatment strategies. G-Transformer supports g-computation, a causal inference technique for estimating treatment effects under dynamic treatment regimes (Robins, 1986, 1987). G-Transfomer estimates the conditional distribution of relevant covariates given covariate and treatment history at each time point using an encoder architecture, then produces Monte Carlo estimates of counterfactual outcomes by simulating forward patient trajectories under treatment strategies of interest. We evaluated G-Transformer extensively using multiple simulated datasets from two mechanistic models where ground-truths under the counterfactual strategies can be measured, and a real-world sepsis dataset to assess G-Transformer’s potential clinical utility in counterfactual predictions under alternative dynamic fluid administration regimes. Our contributions are the following:
  • G-Transformer architecture. We introduce a novel Transformer-based architecture that enables counterfactual predictions under dynamic and time-varying treatment strategies to provide estimates of individual or population-level treatment effects. We present a custom sequential training procedure for supporting g-computation that performs better than the seq2seq training typically used in Transformers.
  • Evaluation under static time-varying treatment regimes. Using a simulated tumor growth dataset, we demonstrated that G-Transformer out-performed other state-of-the-art deep learning approaches, including rMSN (Lim et al., 2018), CRN (Bica et al., 2020a), G-Net (Li et al., 2021), and Causal Transformer (Melnychuk et al., 2022), in counterfactual prediction under static time-varying treatments, in which treatments are time-varying but do not depend on past covariate history.
  • Evaluation under dynamic and time-varying treatment regimes. We used CVSim (Heldt et al., 2010), a mechanistic model of the cardiovascular system, to simulate counterfactual patient trajectories under various dynamic fluid and vasopressor administration strategies, and demonstrated that G-Transformer out-performed other baselines, including G-Net and a linear implementation of g-computation, in counterfactual prediction under dynamic and time-varying treatment regimes.
  • Evaluation under real-world ICU dataset. Using real-world data of a sepsis cohort from the MIMIC-IV database (Johnson et al., 2023b), we evaluated the performance of G-Transformer in predicting outcomes under observational treatment regimes, and demonstrated that G-Transformer can generate clinically meaningful counterfactual predictions under alternative dynamic fluids administration regimes.
Generalizable Insights about Machine Learning in the Context of Healthcare
Our extensive experiments show strong empirical performance of G-Transformer under both static and dynamic time-varying treatment strategies over other state-of-the-art techniques. G-Transformer can facilitate treatment decision making at individual patient level by estimating patient trajectories under different treatment strategies of interest. At a population health level, G-Transformer can potentially be used to estimate population-level treatment effects to inform clinical guidelines and health policies. Although we focus on applications in healthcare in this proposed work, the G-Transformer architecture can be applied to other sequential decision making tasks that involve dynamic and time-varying interventions.
2 Related Work
Recent works by Lim et al. (2018); Bica et al. (2020a, b) presented deep learning approaches to estimate time-varying treatment effects. Bica et al. (2020a) applied ideas from domain adaptation to estimate treatment effects over time while Lim et al. (2018) used RNN regression models with history adjusted marginal structural models (MSMs) (Van der Laan et al., 2005) to make counterfactual predictions. Causal Transformer (CT) has recently been proposed a Transformer-based architecture to estimate counterfactual outcomes (Melnychuk et al., 2022). However, previous approaches focus on estimating counterfactual outcomes under time-varying treatment strategies where treatments are not dependent on past covariate history. None of these approaches are designed to estimate effects under dynamic treatment strategies, in which treatment depends on recent covariate history. In contrast, our work focus on estimating treatment effects under dynamic treatment strategies, in which treatments depend on covariate history.
Dynamic-MSMs can be used to estimate expected counterfactual outcomes under restricted classes of simple dynamic treatment regimes (Orellana et al., 2010; Shahn et al., 2020). It performs time-varying confounding adjustment via inverse probability weighting. The limitation with dynamic-MSMs is that the optimal treatment strategy of realistic complexity may not be included in the restricted class considered.
G-computation can be used to estimate the average effect of a dynamic treatment regime (DTR) on the population, or the conditional effect given observed patient history. Previous implementations of g-computation used classic linear models (Taubman et al., 2009) or recurrent neural networks based architectures, such as LSTM (Li et al., 2021). G-Net (Li et al., 2021) is based on g-computation and used LSTMs for counterfactual prediction of time-varying treatment outcomes under alternative dynamic treatment strategies. In contrast, our work focused on a Transformer-based architecture to better capture complex dependencies in time-varying data. DeepACE (Frauen et al., 2023) is a recurrent neural network based model that leverages the g-computation framework to address time-varying confounding, but primarily focuses on estimating the Average Causal Effect (ACE).
Recent work by Wu et al. (2023) presented an alternating sequential model that used Transformers for clinical treatment outcome prediction, but their work did not support counterfactual prediction. Seedat et al. (2022) and Holt et al. (2024) focused on Continuous-Time Modeling, attempting to predict counterfactual treatment effects over irregular time lengths. They utilized neural controlled differential equations to model patient historical trajectories and attempted to make uncertainty estimates based on this model. T4 (Liu et al., 2023) introduced a LSTM-based neural network architecture for predicting the counterfactual treatment effects of sepsis. They also estimated the uncertainty associated with treatment effects.
3 Methods
3.1 Background: G-computation and Problem Setup
G-computation Robins (1986) can be used to estimate the average effect of a dynamic treatment regime (DTR) on the population, or the conditional effect given observed patient history. In the latter case, we can estimate the expected counterfactual trajectory of outcomes of interest under alternative treatment strategies of interest for a particular patient history.
Problem Setup
Our goal is to predict patient outcomes under alternative future treatment strategies given observed patient histories. We follow notations introduced in Li et al. (2021) in this study.
Let:
  • $t\in\{0,\ldots,K\}$ denote time, assumed discrete, with $K$ being the end of followup;
  • $A_{t}$ denote the observed treatment action at time $t$;
  • $Y_{t}$ denote the observed outcome at time $t$
  • $L_{t}$ denote a vector of covariates at time $t$ that may influence treatment decisions or be associated with the outcome;
  • $\bar{X}_{t}$ denote the history $X_{0},\ldots,X_{t}$ and $\underline{X}_{t}$ denote the future $X_{t},\ldots,X_{K}$ for arbitrary time varying variable $X$.
At each time point, we assume the causal ordering $(L_{t},A_{t},Y_{t})$, i.e. the treatments are administered after having observed the covariates at time step $t$. Let $H_{t}\equiv(\bar{L}_{t},\bar{A}_{t-1})$ denote patient history preceding treatment at time $t$. A dynamic treatment strategy $g$ is a collection of functions $\{g_{0},\ldots,g_{K}\}$, one per time point, such that $g_{t}$ maps $H_{t}$ onto a treatment action at time $t$. An example dynamic strategy might be to administer treatment A if the patient’s vital signs, e.g. arterial blood pressure, drops below certain threshold.
Let $Y_{t}(g)$ denote the counterfactual outcome that would be observed at time $t$, had treatment strategy $g$ been followed from baseline (Robins, 1986). Further, let $Y_{t}(\bar{A}_{m-1},\underline{g}_{m})$ with $t\geq m$ denote the counterfactual outcome that would be observed had the patient received their observed treatments $\bar{A}_{m-1}$ through time $m-1$ then followed strategy $g$ from time $m$ onward. Here, the treatment strategy $g$ is typically specified by the domain experts, e.g. clinicians. In this work, the outcome $Y_{t}$ can be deemed to be a variable in the vector $L_{t+1}$.
In counterfactual point prediction, our goal is to estimate expected counterfactual patient outcome trajectories
$\{E[Y_{t}(\bar{A}_{m-1},\underline{g}_{m})|H_{m}],t\geq m\}$ (1)
given observed patient history through time $m$ for any $m$ and any specified treatment strategy $g$, where $g$ is specified by a domain expert, e.g. a clinician. It may be of interest to estimate the counterfactual outcome distributions at future time points
$\{p(Y_{t}(\bar{A}_{m-1},\underline{g}_{m})|H_{m}),t\geq m\}.$ (2)
The g-computation algorithm enables the estimation of (1) and (2) under the following assumptions (Robins, 1986): Consistency, Sequential Exchangeability, and Positivity.
Assumption 3.1. Consistency asserts that the observed outcome aligns with the counterfactual outcome that would have occurred had the observed treatment been applied.
Assumption 3.2. Sequential exchangeability requires that there is no unobserved confounding, i.e., all drivers of treatment decisions prognostic for the outcome are included in the data.
Assumption 3.3. Positivity posits that the counterfactual treatment strategy of interest has some non-zero probability of actually being implemented.
Under assumptions 3.1-3.3, for $t=m$ we have that
$p(Y_{m}(\bar{A}_{m-1},g_{m})|H_{m})=p(Y_{m}|H_{m},A_{m}=g_{m}(H_{m})),$ (3)
i.e. the conditional distribution of the counterfactual is the conditional distribution of the observed outcome given patient history and given that treatment follows the strategy of interest. For $t>m$, we need to adjust for time-varying confounding. With $X_{i:j}=X_{i},\ldots,X_{j}$ for any random variable X, under assumptions 1-3 the g-formula yields
$p(Y_{t}(\bar{A}_{m-1},\underline{g}_{m})=y|H_{m})$
$=\int_{l_{m+1:t}}p(Y_{t}=y|H_{m},L_{m+1:t}=l_{m+1:t},A_{m:t}=g(H_{m:t}))$
×$\quad\quad\quad\quad\times\prod_{j=m+1}^{t}p(L_{j}=l_{j}|H_{m},L_{m+1:j-1}=l_{m+1:j-1},$
$\quad\quad\quad\quad\quad\quad\quad\quad A_{m,j-1}=g(H_{m},l_{m+1:j-1})).$ (4)
We can approximate this integral through Monte-Carlo simulation.
G-computation algorithm requires the ability to simulate from joint conditional distributions $p(L_{t}|\bar{L}_{t-1},\bar{A}_{t-1})$ of the covariates given patient history at time $t$. These conditional distributions need to be estimated from data. Most implementations of g-computation use generalized linear regression models (GLMs) to estimate the conditional distributions of the covariates. These models often do not capture the complex temporal dependencies in the patient data. We propose the G-Transformer for this task.
3.2 G-Transformer Architecture
Figure 1: G-Transformer Architecture. G-Transformer uses two distinct encoders for categorical ($L^{ca}$) and continuous variables ($L^{co}$) respectively. The diagram illustrates (i) the training process through time steps T-1 and T, and (ii) the counterfactual simulation process starting at time T, with the counterfactual strategy initiated at the end of time T-1. Teacher forcing is used during training.
We utilized two Transformer encoders as the sequential model to separately learn hidden representations for continuous and categorical covariates in G-Transformer. We use the teacher forcing in the training process and g-computation at simulation process. The architecture is shown in Fig.1.
Let $L_{t}^{0},...,L_{t}^{p}$ denote $p$ components of the vector of covariates $L_{t}$. When we calculate joint conditional distributions $p(L_{t}|\bar{L}_{t-1},\bar{A}_{t-1})$ at time t, we leverage the conditional probability identity as in (Li et al., 2021):
$p(L_{t}|\bar{L}_{t-1},\bar{A}_{t-1})$ $=p(L^{0}_{t}|\bar{L}_{t-1},\bar{A}_{t-1})$
×$\times p(L^{1}_{t}|L^{0}_{t},\bar{L}_{t-1},\bar{A}_{t-1})$
[…]
×$\times p(L^{P-1}_{t}|L^{0}_{t},\ldots,L^{P-2}_{t},\bar{L}_{t-1},\bar{A}_{t-1})$ (5)
In G-Transformer, we divide covariates into two groups: one for categorical variables $L_{t}^{ca}$, and the other for continuous variables $L_{t}^{co}$. Thus, $p(L_{t}|\bar{L}_{t-1},\bar{A}_{t-1})$ is defined as:
$p(L_{t}|\bar{L}_{t-1},\bar{A}_{t-1})=p(L^{ca}_{t}|\bar{L}_{t-1},\bar{A}_{t-1})\times p(L^{co}_{t}|L^{ca}_{t},\bar{L}_{t-1},\bar{A}_{t-1})$ (6)
Although assigning a customized model to each covariate may achieve better performance, considering the limitations of computational resources and the potential dependencies among variables, we decide to have only two Transformer encoder models to perform representation learning for categorical variables and continuous variables, respectively. At each time step $t$, we can compute represenetation of patient history $R_{t}=r_{t}(\bar{L}_{t},\bar{A}_{t};\theta)$, where $\theta$ denotes the learnable parameters from model $r_{t}$. We will introduce the specific content of model $r_{t}$ in the following section.
Representation Learning. For the input of categorical encoder, we concatenate all the historical treatments $\bar{A}_{t}$ and covariates $\bar{L}_{t}$ together, and use a linear layer to obtain their combined representation $h_{ca}^{r}$:
$h_{ca}^{r}=Linear(concat(\bar{A}_{t},\bar{L}_{t}))$ (7)
Afterwards, we feed the combined representations of treatments and outcomes $h_{ca}^{e}$ into a Transformer encoder to obtain the final representation of past covariate history:
$h_{ca}=Transformer\>Encoder_{ca}(h_{ca}^{r})$ (8)
The internal architecture of the Transformer encoder sublayers follows the original Transformer paper (Vaswani et al., 2017), including a multi-head self-attention layer, feed-forward network, and layer normalization mechanism. After obtaining $h_{ca}$, we map the hidden representation to the predicted future time step through a linear layer:
$\bar{L}_{t+1}^{ca}=Linear(ReLU(h_{ca}))$ (9)
During the training phase, we use teacher forcing to guide the training of the categorical and continuous encoders. In addition to the patient’s historical sequences $(\bar{A}_{t},\bar{L}_{t})$, we also use observed value $L_{t+1}^{ca}$ as input into the continuous encoder:
$h_{co}^{r}=Linear(concat(\bar{A}_{t},\bar{L}_{t},L_{t+1}^{ca})).$ (10)
As with categorical variables, we use a Transformer encoder to obtain a hidden representation $h_{co}$ that incorporates the guiding variables. Then we utilize a linear layer to modify the hidden state to final output $\bar{L}_{t+1}^{co}$:
$h_{co}=T$ $ransformer\>Encoder_{co}(h_{co}^{r})$ (11)
$\bar{L}_{t+1}^{co}=Linear(h_{co})$ (12)
3.3 Training of G-Transformer
Since G-computation attempts to model the joint conditional probability distribution of the next time step $p(L_{t}|\bar{L}_{t-1},\bar{A}_{t-1})$ given the observable sequence, this approach is more aligned with the modeling methods of some RNN-like linear regression models. However, due to the Transformer’s nature as a seq-2-seq model, its multi-head self-attention mechanism endows it with powerful parallel computing capabilities. Therefore, the Transformer encoder is more inclined to model the probability distribution of a sequence, that is $p(\bar{L}_{2:t}|\bar{L}_{1:t-1},\bar{A}_{1:t-1})$. We tried two different training methods on our counterfactual prediction task, and the results are shown in the Appendix A.1. Furthermore, in real-world scenarios, most treatment variables appear in binary form, with few treatment variables presented in continuous form. Therefore, we believe that prioritizing modeling for categorical variables can provide more insights for the prediction of continuous variables. This approach also aligns more closely with the fundamental procedural form we use in the simulation process to calculate treatment variable $a_{m}=g_{m}(H_{m})$ in Algorithm. 2. After completing the entire prediction process, we use cross-entropy loss and MSE (Mean Squared Error) loss to optimize the prediction results for categorical and continuous variables, respectively. For categorical variables:
$\mathcal{L}_{ce}=-\frac{1}{N\times P_{ca}}\sum_{i}^{N}\sum_{c=1}^{P_{ca}}y_{ic}\;log(p_{ic};\theta)$ (13)
where $P_{ca}$ denotes the number of categorical variables, N denotes the number of paitents, and $\theta$ denotes the model parameters. In fact, we separately calculated the cross-entropy losses of different variables and computed their averages to account for the needs of multi-class variables. For continuous variables, we utilize the MSE loss:
$\mathcal{L}_{mse}=\frac{1}{N}\sum_{i=1}^{N}\sum_{t=1}^{K}||\bar{L}_{t,i}^{co}-\hat{L}_{t,i}^{co}||^{2}$ (14)
where $\hat{L}_{t,i}^{co}$ denotes the output of continuous encoder part of G-Transformer for patient $i$ in time $t$, $N$ denotes number of patients and $K$ denotes length of time. Algorithm 1 illustrates the overall training process of G-Transformer.
Input: Training set $\mathcal{D}$ data tuples: $\mathcal{D}=\{(\bar{L},\bar{A})\}$ ; the number of the prediction length $l$, Model parameters $\theta_{ca}$ and $\theta_{co}$; the number of the learning epochs $E$.
while $e<E$ do
       for $t=2$ to $l$ do
             Concatenate treatment variables and outcome variables $\bar{Y}_{1:t-1}=\{\bar{L}_{1:t-1},\bar{A}_{1:t-1}\}$.
             Use categorical encoder to compute the $\bar{Y}_{2:t}^{ca}=f_{ca}(\bar{Y}_{1:t-1};\theta_{ca})$.
             Keep the last time step of the predicted categorical variable $\bar{Y}_{t}^{ca}=\{\bar{L}_{t}^{ca},A_{t}^{ca}\}$.
             Use teacher forcing to process the input of continuous encoder $\bar{Y}_{1:t-1}^{te}=\{\bar{L}_{1:t-1},\bar{A}_{1:t-1},\bar{L}_{t}^{ca}\}$.
             Use continuous encoder to compute the $\bar{Y}_{2:t}^{co}=f_{co}(\bar{Y}_{1:t-1}^{te};\theta_{co})$.
             Keep the last time step of the predicted continuous variable $\bar{Y}_{t}^{co}=\{\bar{L}_{t}^{co},A_{t}^{co}\}$.
             Save the last time step $Y_{t}^{co},Y_{t}^{ca}$.
       end for
      Concatenate the results of all predicted time steps $\bar{Y}_{2:t}^{ca}=\{{Y}_{2}^{ca},...,{Y}_{t}^{ca}\}$ and $\bar{Y}_{2:t}^{co}=\{{Y}_{2}^{co},...,{Y}_{t}^{co}\}$.
       Update model parameters $\theta_{ca}$ and $\theta_{co}$ by minimizing (13) and (14) with gradient descent
end while
return updated model parameters $\theta_{ca}$ and $\theta_{co}$
Algorithm 1 G-Transformer training algorithm
3.4 Simulation
After having a trained sequential model $f(R;\theta)$, we can obtain the conditional expectation estimate $\mathbb{E}[L_{t}|\bar{L}_{t-1},\{L_{t}^{i}\}_{i=1,...,p},\bar{A}_{t-1}]$ of all $p$ covariates at time step $t$ through Monte Carlo simulation. Without making parametric assumptions, we could simulate from:
$L_{t}^{p}\simeq\mathbb{\hat{E}}[(L_{t}^{p}|L_{t}^{0},...,L_{t}^{p-1},\bar{L}_{t-1},\bar{A}_{t-1})]+\epsilon_{t}^{p}$ (15)
where $\epsilon_{t}^{p}$ is a noise which draws from the empirical distribution of the residuals $L^{p}_{t}-\hat{L}^{p}_{t}$ in a holdout set. If the variable is binary, then $\epsilon_{t}^{p}$ here is sampled from a Bernoulli distribution parameterized by empirical data. Afterward, we use Algorithm 2 for $K$ times to perform multi-timestep simulations. Figure 1 illustrates the simulation process. Different from the training process, since we cannot access to the ground truth, we sample from the learned conditional distribution when performing counterfactual predictions. For categorical variables, we use the parameters derived from the result $\hat{L}^{ca}$ of the categorical encoder to parameterize a categorical distribution (or a Bernoulli for binary indicator variables), and sample from the corresponding distribution. For the continuous variables, we added noise sampled from the empirical distribution collected from the validation set to the output part of the continuous encoder.
Set $a_{m}^{*}=g_{m}(H_{m})$
Sample $l_{m+1}^{*}$ from $p(L_{m+1}|H_{m},A_{m}=a_{m}^{*})$
Set $a_{m+1}^{*}=g_{m}(H_{m},l_{m+1}^{*},a_{m}^{*})$
Sample $l_{m+2}^{*}$ from $p(L_{m+2}|H_{m},L_{m+1}=l_{m+1}^{*},A_{m}=a_{m}^{*},A_{m+1}=a_{m+1}^{*})$
Continue simulations through time $K$
Algorithm 2 Simulation (one time-step)
We repeat the g-computation simulation Algorithm $M$ times. At the end of this process, we have $M$ simulated draws of the counterfactual outcome for each time $t=\{m,\ldots,K\}$. For each $t$, the empirical distribution of these draws constitutes a Monte-Carlo approximation of the counterfactual outcome distribution (2). The sample averages of the draws at each time $t$ are an estimate of the conditional expectations (1) and can serve as point predictions for $Y_{t}(\bar{A}_{m-1},\underline{g}_{m})$ in a patient with history $H_{m}$.
3.5 Evaluation
To evaluate G-Transformer, we use simulated data in which counterfactual ground truth can be known. Specifically, we use CVSim (Heldt et al., 2010), a well-established mechanistic model of the cardiovascular system, to simulate counterfactual patient trajectories under various dynamic fluid and vasopressor administration strategies. Additionally, using a simulated tumor growth data set, we compare G-Transformer with recently introduced G-Net (Li et al., 2021), Recurrent Neural Networks (CRN) (Bica et al., 2020a), a Recurrent Marginal Structural Network (R-MSN) (Lim et al., 2018), a recurrent neural network implementation of a history adjusted marginal structural model for estimating static time-varying treatment effects, and Causal Transformer (Melnychuk et al., 2022).
We applied our proposed G-Transformer approach to predicting outcomes of sepsis patients in the ICU under alternative fluid resuscitation treatment regimes (e.g. aggressive vs. conservative) using de-identified real-world intensive care units (ICU) data from the MIMIC-IV (Johnson et al., 2023b) publicly available from PhysioNet.
We quantitatively evaluate G-Transformer’s performance in predicting patient trajectories under observational treatment regimes learned from the ICU data. This entails using the G-Transformer architecture to predict treatment at each time-step under the ”usual care” in the observational data. Under predictive check, the treatments are from G-Transformer predicted actions conditioned on patient history.
4 Experiments on Simulated Data from Mechanistic Models
4.1 Cancer Growth Experiments and Results
Cancer Growth Data Generation
As in Lim et al. (2018); Bica et al. (2020a); Li et al. (2021); Melnychuk et al. (2022), we generate simulated ‘observational’ data from a pharmacokinetic-pharmacodynamic model of tumor growth under a stochastic regime (Geng et al., 2017). In this simulation, chemotherapy and radiation therapy comprise a two dimensional time-varying treatment impacting tumor growth. Under the observational regime, probability of receiving each treatment at each time depends on volume history, so there is time-varying confounding. We use the same experiment settings and data generation procedure as reported in Li et al. (2021). We generated four test sets in which four counterfactual regimes were followed for the final four time points in the test set: (1) give only radiotherapy, (2) only chemotherapy, (3) both chemotherapy and radiotherapy, and (4) no treatment. These counterfactual regimes are static time-varying regimes (in which treatments do not depend on prior patient history), as rMSN, CRN, and Causal Transformer are intended to estimate counterfactuals under static time-varying regimes.
Table 1 presents the percent root mean square error (RMSE) of predictions from rMSN, CRN, linear implementation of g-computation, G-Net, Causal Transformer, and G-Transformer in the final four time steps when counterfactual strategies were in effect, conditioned upon previous time points. The raw RMSE values were divided by the maximum possible tumor volume, 1150 $cm^{2}$, as in Lim et al. (2018), to calculate the percent RMSE.
We observe that G-Transformer achieved the best overall RMSE performance in three of the four cancer growth datasets, demonstrating its advantages in counterfactual prediction task compared to other state-of-the-art deep learning models, including rMSN, CRN, G-Net and Causal Transformer. Causal Transformer outperformed other models under the chemotherapy (Chemo) counterfactual regime, but fall short in all other counterfactual regimes.
Table 1: Cancer growth data: Percent RMSE for various prediction horizons. A = Overall. CT = Causal Transformer. Best performing models in bold.
rMSN CRN Linear G-Net CT Ours
$t$ g-comp G-Transformer
No 1 1.13 1.00 0.63 0.25 0.67 0.26
Treat 2 1.24 1.20 1.21 0.47 0.65 0.44
3 1.85 1.49 1.78 0.72 0.63 0.55
4 2.60 1.78 2.35 1.01 0.63 0.66
A 1.68 1.40 1.62 0.67 0.65 0.48
Radio 1 5.27 4.91 7.14 3.29 5.15 2.74
2 5.38 4.92 7.43 3.14 2.98 2.50
3 5.13 4.94 7.05 3.02 2.72 2.40
4 4.81 4.92 6.50 3.00 2.53 2.52
A 5.15 4.92 7.04 3.11 3.35 2.54
Chemo 1 1.42 1.04 1.58 0.34 0.66 0.36
2 1.27 1.09 3.14 0.63 0.50 1.16
3 1.46 1.03 4.52 0.84 0.44 1.95
4 1.69 1.02 5.47 0.89 0.41 2.45
A 1.47 1.05 3.96 0.71 0.50 1.48
Radio 1 4.76 4.66 7.76 3.10 8.40 2.48
Chemo 2 3.59 4.36 7.32 2.28 7.72 2.04
3 2.76 3.65 6.13 1.50 7.79 1.54
4 2.30 2.95 4.88 1.18 7.67 1.10
A 3.48 3.96 6.62 2.15 7.90 1.79
4.2 CVSim Experiments and Results
CVSim Data Generation
To evaluate counterfactual predictions, it is necessary to use simulated data in which counterfactual ground truth for outcomes under alternative treatment strategies is known. To this end, we performed experiments on data generated by CVSim, a program that simulates the dynamics of the human cardiovascular system (Heldt et al., 2010). We used a CVSim 6-compartment circulatory model which takes as input 28 variables that together govern a hemodynamic system. We built on CVSim by adding stochastic components and interventions for the purposes of evaluating our counterfactual simulators.
We generated an ‘observational’ dataset $D_{o}$ under treatment regime $g_{o}$ and two ‘counterfactual’ datasets $D_{c1}$ and $D_{c2}$ under treatment regimes $g_{c1}$ and $g_{c2}$. The data generating processes producing $D_{o}$ and $D_{cj}$ were the same except for the treatment assignment rules. For each $j$, $g_{cj}$ was identical to $g_{o}$ for the first $m-1$ simulation time steps before diverging to a different treatment rule for time steps $m$ to $K$. Full details are in the Appendix C.
Experiment
As in Li et al. (2021), we generated a total of 12,000 trajectories in $D_{o}$ ($N_{o}=12,000$), of which 80% were used for training, and the remaining 20% for validation. For testing, we generated 1000 observations in the $D_{cj}$ datasets ($N_{c}=1000$). We included a total of 20 output variables (i.e. two treatment variables and 18 covariates influencing treatment assignment under $g_{o}$) from CVSim to construct $D_{o}$ and $D_{cj}$; each trajectory is of length 66 time steps (d=20, K=66). In each $D_{cj}$, the switching time point $m$ from $g_{o}$ to $g_{c}$ is fixed at 34 for all trajectories ($m=34$).
Given observed covariate history through 34 time steps and treatment history through 33 time steps of each trajectory in each $D_{cj}$, we computed both the population-level and individual-level RMSE, as well as the calibration, of counterfactual predictions from G-Transformer and its counterparts for time steps 35 to 66. Moreover, we presented the estimated and actual population-level average trajectories under gc1 and gc2 for selected variables. We also visually demonstrated the treatment effects estimation.
Population-Level and Individual-Level RMSE
We use both population-level and individual-level RMSE to assess the predictive accuracy of the model for continuous variables in CVSim data. To derive population-level RMSE, we aggregated the simulated trajectories from all patients into a single population-level trajectory. This predicted average population-level trajectory was then compared to the ground-truth average population-level trajectory for RMSE calculation. Individual-level RMSE is calculated by first determining the square root of the mean squared differences between each patient’s observed and predicted trajectories. For G-Transformer, G-Net and the linear implementation of g-computation, the predicted trajectory is averaged from 100 Monte Carlo simulations conditioned on historical data up to time (k-1). The RMSE values across all patients are then averaged to derive individual-level RMSE.
Table 2 presents a comparison of individual-level and population-level RMSE among linear implementation of g-computation, G-Net, Causal Transformer, and G-Transformer under gc1 and gc2. The G-Transformer outperforms linear implementation of g-computation and G-Net, and considerably outperforms Causal Transformer in this counterfactual prediction task. It should be noted that Causal Transformer Melnychuk et al. (2022) can estimate counterfactual outcomes under time-varying treatments, but it is not designed to estimate counterfactual outcomes under dynamic treatment regimes, where treatments depend on past covariate history. Nevertheless, we present its performance here as a baseline.
Table 2: CVSim: Counterfactual prediction under gc1 and gc2. Performance of various models in predicting remaining 32-hour patient covariate trajectories based on first 34-hour patient covariate trajectories. Values reported represent the population-level and individual-level RMSE between the predicted and counterfactual trajectories.
Population Level Individual Level
Model RMSEgc1 RMSEgc2 RMSEgc1 RMSEgc2
Linear(g-comp) 0.252 0.272 1.103 1.193
G-Net 0.061 0.105 1.020 1.114
Causal Transformer 0.885 1.099 1.435 1.727
G-Transformer 0.050 0.054 1.016 1.104
Calibration
We assess the calibration of a G-Transformer, G-Net, or linear implementation of g-computation, denoted as $G$, as follows. Given lower and upper quantiles $\alpha_{low}$ and $\alpha_{high}$, the calibration measures the frequency with which the actual counterfactual covariate $L_{ti}^{h,cj}$ is between the $\alpha_{low}$ and $\alpha_{high}$ quantiles of the $M$ simulations $\{\tilde{L}_{ti}^{h,cj}(H_{mi}^{cj},G,k):k\in 1:M\}$. If this frequency is approximately $\alpha_{high}-\alpha_{low}$, then $G$ is well calibrated.
Figure 2 depicts calibration for G-Transformer, G-Net, and linear implementation of g-computation with lower and upper quantiles of 0.05 and 0.95. G-Transformer generally performed better than G-Net and linear implementation of g-computation over all time steps. The model is well calibrated at the beginning, but as simulation time steps increases, calibration becomes weaker for all models. This is in line with the RMSE increase over time as can be seen from Table 7 in the Appendix A.3.
(a) Calibration under $gc_{1}$
(b) Calibration under $gc_{2}$
Figure 2: Calibration over time for $g_{c1}$ and $g_{c2}$
Population-Level Trajectories
The G-Transformer, G-Net, and linear implementation of g-computation can be used to quantitatively demonstrate counterfactual outcomes through population average trajectories. The estimated and actual population-level average trajectories under gc1 and gc2 for selected variables, specifically arterial flow (AQ), arterial pressure (AP), and heart rate (HR), are presented in Figure 3 (3(a)) to (3(f)). The plots show that G-Transformer outperforms G-Net and linear implementation of g-computation, and more accurately predict AQ, AP, and HR counterfactual trajectories from the population level.
(a) AQ ($gc_{1}$)
(b) AP ($gc_{1}$)
(c) HR ($gc_{1}$)
(d) AQ ($gc_{2}$)
(e) AP ($gc_{2}$)
(f) HR ($gc_{2}$)
(g) AQ ($gc_{1}$-$gc_{2}$)
(h) AP ($gc_{1}$-$gc_{2}$)
(i) HR ($gc_{1}$-$gc_{2}$)
Figure 3: Plots (3(a)) to (3(f)) show the estimated and actual population average trajectories for three selected variables: arterial flow (AQ), arterial pressure (AP), and heart rate (HR), according to three models’ predictions under gc1 vs gc2. Plots (3(g)) to (3(i)) show population-level treatment effect of $g_{c1}$ vs $g_{c2}$ estimated by G-Transformer vs G-Net.
Treatment Effects
The G-Transformer, G-Net, and linear g-computation implementation can also be used to estimate population average counterfactual outcomes and treatment effects, quantities sometimes more relevant to policy decisions than individual-level counterfactual predictions. Figure 3 (3(g)) to (3(i)) show the estimates and true values of the population average treatment effect of following $g_{c1}$ as opposed to $g_{c2}$ on selected variables. We see that the G-Transformer effectively estimates these population-level quantities of interest.
5 Experiments on MIMIC Data
5.1 MIMIC Cohort
The data employed in this study was extracted from the Medical Information Mart for Intensive Care IV database (MIMIC-IV v1.0), containing medical records from more than 523,500 hospital admissions and 76,500 ICU stays at the Beth Israel Deaconess Medical Center (BIDMC) between 2008 and 2019 Johnson et al. (2023a). Our cohort consists of ICU patient identified as septic under the Third International Consensus Definitions for Sepsis and Septic Shock (Sepsis-3) Singer et al. (2016) and did not meet exclusion criteria (see Appendix D for details). Our final sepsis cohort consisted of 8,934 total patients which we further split into 7,147 patients, 893 patients, and 894 patients for the training, validation, and test data respectively.
5.2 Study Design
Since there is no ground-truth observations available for the MIMIC data for counterfactual predictions, we first assess our models’ performance through a “predictive check.” This involved conducting Monte Carlo simulations and projecting forward patient covariate trajectories within the test set under the observational regime. Subsequently, we compared these simulated trajectories (averaged across 100 Monte Carlo simulations per patient) to the ground-truth data. Following predictive check, we assess our model in performing counterfactual predictions.
Predictive Checks
We compared the 24-hour trajectories between predicted and ground-truth under G-Transformer, G-Net, and Causal Transformer, using individual-level and population-level RMSE as metrics. The method for calculating the population-level RMSE is as described in the previous section. However, since patient trajectories can be censored due to events such as death and release in MIMIC data, individual-level RMSE is adjusted to account for error due to over- and under-prediction in the trajectory length as in Su et al. (see Appendix B.1 for details).
We also compared the models’ performance in predicting the occurrence of adverse clinical outcomes during the first 24-hours patients were in the ICU. Confidence intervals around AUCs are generated based on techniques described in DeLong et al. (1988). To generate a probability for the 24-hour window, we first assigned a binary label for each Monte Carlo simulation based on whether the outcome is predicted to have occurred: $1$ if outcome of interest was predicted to occur between timestamps $k$ to $24$, and 00 otherwise. Then the average across the Monte Carlo simulations per patient was used as the predicted probability for that patient. The ground truth was determined in the same manner by looking for the presence of the outcome of interest in the ground-truth trajectory.
Counterfactual Experiments
We adapted our counterfactual strategies from established clinical trials studying the early treatment of sepsis. This strategy imposes a fluid cap on the total amount of fluids intake of a patient to $X$ liters, and ceases administration once the fluid cap is reached or when the patient is fluid-overloaded. More specifically, this strategy was based on the Crystalloid Liberal or Vasopressors Early Resuscitation in Sepsis (CLOVERS) clinical trial (Self et al., ). For a patient with blood pressure below 65mmHg at time $t$, a 1000mL bolus was administered if the total volume of fluids (including both treatment and maintenance) they received up until that time point did not exceed $X$ liters and if they did not exhibit any signs of fluid overload (as indicated by the presence of pulmonary edema based on chest x-ray radiology reports). We experimented with imposing a fluid cap of $3$L and $5$L as fluid conservative strategies, and a fluid strategy with no fluid cap to represent a fluid liberal strategy.
5.3 Predictive Check: Population-Level and Individual-Level RMSE
Table 3 presents the predictive check on population-level and individual-level RMSE for continuous variables. We aim to assess the performance of G-Transformer, G-Net, and Causal Transformer in predicting 24-hour patient covariate trajectories, starting at time $k$ conditioned on covariates from the previous $k-1$ timesteps of post-ICU admission. $k$ ranges from 2 to 8 hours after ICU admission. Values reported represent the RMSE of continuous covariates between the predicted and ground-truth trajectories. To account for mismatch in trajectory length due to over- and under-prediction of death and release time, for both predicted and actual trajectories, depending on nature of covariates, post-death timesteps are padded with one of the following values: normalized zero, normalized minimum value in dataset, or normalized logarithmic zero. The post-release timesteps are consistently padded with normalized population mean.
According to Table 3, we can see that errors in predicted 24-hour patient trajectories are relatively stable for G-Transformer and Causal Transformer, and G-Transformer consistently outperforms G-Net and Causal Transformer across different values of $k$. Although Causal Transformer Melnychuk et al. (2022) is not intended to estimate counterfactual outcomes under dynamic treatment regimes (where treatments depend on past covariate history), nevertheless, we present its performance here as a baseline.
Table 3: MIMIC-IV Experiments: Predictive check on continuous variables at population-level and individual-level RMSE. $k$ is the simulation start time relative to the ICU admission time. Performance in predicting 24-hour patient covariate trajectories, starting at hour $k$ conditioned on covariates from the previous $k-1$ timesteps of post-ICU admission. $k$ ranges from 2 to 8 hours after ICU admission. Values reported represent the population-level and individual-level RMSE of continuous covariates between the predicted and ground-truth trajectories. Individual-level RMSE is adjusted to account for error due to over- and under-prediction in the trajectory length as a result of death and discharge.
Population Level RMSE Individual Level RMSE
Model / k (hours) 2 4 6 8 2 4 6 8
G-Net 0.338 0.265 0.214 0.181 15.85 13.71 11.40 9.29
Causal Transformer 2.770 2.736 2.643 2.514 14.37 15.00 15.73 16.54
G-Transformer 0.297 0.233 0.204 0.178 7.96 8.83 9.46 8.62
Predictive Check: an Illustrative Example from G-Transformer’s Monte Carlo Simulations
Figure 4 illustrates the Monte Carlo simulations generated by G-Transformer for selected variables, specifically systolic blood pressure (SBP) and heart rate (HR), given an example MIMIC patient from test dataset. We observe that the predicted trajectory averaged from 100 Monte Carlo simulations follow closely with the ground-truth trajectory.
(a) SBP
(b) HR
Figure 4: Illustration of predictive check on an example MIMIC test set patient. G-Transformer simulated systolic blood pressure (SBP) and heart rate (HR) trajectories (100 Monte Carlo simulations in light blue, average in solid dark blue) compared with ground truth (red) for one patient under predictive check (simulation starting at hour 6).
Predictive Check: 24-Hour Outcomes Predictions
Table 4 presents AUCs (and 95% CI) in predicting clinical outcomes within the first 24 hours of ICU admission at varying simulation start times from G-Transformer and G-Net. There are six outcomes of interest: diagnosis of pulmonary edema, use of a mechanical ventilator (MV), administration of diuretics, dialysis, ICU release, and in-hospital mortality. We observe that the predicted 24-hour clinical outcomes of interest tend to be more accurate as the value of $k$ increases, and G-Transformer generally outperforms G-Net across different values of $k$.
Table 4: Predictive Check. Performance of G-Transformer and G-Net in predicting clinical outcomes within the first 24 hours of ICU admission at varying simulation start times. Values reported are AUCs and 95% confidence intervals.
Model k Edema MV Diuretics Dialysis Release Death
G-Net 2 0.82 (0.79, 0.86) 0.89 (0.87, 0.91) 0.68 (0.63, 0.73) 0.85 (0.75, 0.95) 0.67 (0.61, 0.73) 0.79 (0.67, 0.90)
G-Transformer 2 0.84 (0.80, 0.87) 0.89 (0.86, 0.91) 0.73 (0.68, 0.78) 0.90 (0.81, 0.98) 0.67 (0.61, 0.73) 0.82 (0.71, 0.93)
G-Net 4 0.86 (0.83, 0.89) 0.94 (0.92, 0.95) 0.70 (0.64, 0.75) 0.94 (0.88, 1.00) 0.65 (0.59, 0.71) 0.86 (0.76, 0.96)
G-Transformer 4 0.86 (0.83, 0.89) 0.94 (0.92, 0.96) 0.75 (0.70, 0.80) 0.96 (0.91, 1.00) 0.68 (0.62, 0.74) 0.91 (0.83, 1.00)
G-Net 6 0.89 (0.86, 0.92) 0.95 (0.93, 0.96) 0.73 (0.67, 0.78) 0.93 (0.86, 1.00) 0.68 (0.63, 0.74) 0.83 (0.72, 0.94)
G-Transformer 6 0.87 (0.84, 0.90) 0.94 (0.92, 0.96) 0.74 (0.69, 0.80) 0.96 (0.90, 1.00) 0.69 (0.63, 0.74) 0.89 (0.80, 0.98)
G-Net 8 0.89 (0.87, 0.92) 0.95 (0.94, 0.97) 0.73 (0.68, 0.79) 0.97 (0.92, 1.00) 0.71 (0.66, 0.77) 0.89 (0.79, 0.98)
G-Transformer 8 0.89 (0.86, 0.91) 0.95 (0.93, 0.97) 0.75 (0.70, 0.81) 0.95 (0.89, 1.00) 0.71 (0.66, 0.77) 0.95 (0.79, 0.98)
Counterfactual Prediction: Population-Level Trajectories
We used G-Transformer to simulate covariate trajectories for patients in the test dataset under two conservative fluids strategies (with 3L and 5L fluid cap) and a fluid liberal strategy (with no cap). Selected covariates are presented in Figure 5, allowing for comparison of the three counterfactual regimes. Population-level average trajectories were calculated from the test set (N=894). Counterfactual strategies are applied starting at time step 1, conditioned on observations up until the first hour in the ICU. For each patient, we perform 100 Monte Carlo simulations. The plotted values are averaged across simulated trajectories for all test patients.
Our aim is to assess whether G-Transformer can make logical and clinically meaningful counterfactual predictions. The trends depicted in Figure 5 align with our physiological expectations. Under fluid liberal strategies, the bolus volume, urine output, respiratory rate, and blood pressure exhibit higher values, following the expected trends at a population level. Conversely, the higher levels in creatinine and blood urea nitrogen (BUN) noted under the fluids conservative regime also align with expected outcomes.
Readers should exercise caution when interpreting these plots in Figure 5 as indicators of treatment effects, as the number of patients (or trajectories) contributing to the average may not remain constant throughout the entire 24-hour period. Simulated trajectories might terminate prematurely within the 24-hour period due to death or discharge.
Figure 5: Population-level trajectories under counterfactual fluids strategies. G-Transformer’s predictions for selected covariates under various counterfactual fluids strategies, including two fluids limiting strategies with 3L (blue) and 5L (purple) fluids cap, and a fluids liberal (green) strategy without fluid cap.
6 Discussion and Conclusion
In this paper, we present G-Transformer, a Transformer-based framework supporting g-computation, to estimate clinical outcomes under counterfactual treatment strategies from both synthetic dataset and a real-world sepsis patient cohort. Our experiments using Cancer Growth and CVSim data, demonstrated that G-Transformer achieved superior performance over other state-of-the art approaches. Using a real-world sepsis dataset, we demonstrated that G-Transformer provides clinically meaningful forecasts of covariate trajectories under alternative counterfactual fluid limiting regimes. This study highlights G-Transformer’s potential clinical utility in supporting treatment decisions for ICU patients. Although we focus on applications in healthcare in this study, the G-Transformer architecture can be applied to other sequential decision making tasks that involve dynamic and time-varying interventions.
Limitations
One limitation of this work is that counterfactual predictive density estimates in our experiments do not take into account uncertainty about model parameter estimate. Specifically, given G-Transformer parameters, the distribution of the Monte Carlo simulations produced by g-computation algorithm constitute an estimate of uncertainty about a counterfactual prediction. However, this estimate ignores uncertainty about the G-Transformer parameter estimates themselves. An important area of future work for G-Transformer is adding support for quantification of model uncertainty.
Appendix A Additional CVSim Results
A.1 Alternative G-Transformer Architecture Experiment
Our proposed G-Transformer differs from the traditional Transformer, including the Causal Transformer, in the way it is iteratively trained on time-series data. We concatenated the results of the last time step of each sub-sequence to form a complete prediction sequence for optimization. In contrast, traditional Transformers directly use a sequence-to-sequence framework, shifting the training sequence to obtain the sequence that needs to be predicted. We compared the performance of models implemented with these two training methods on the CVSim dataset, and the results are shown in Table 5.
We can observe that models trained using the sequence-to-sequence approach, including the sequence-to-sequence G-Transformer (Seq2seq G-Transformer) and Causal Transformer, perform notably worse in term of individual-level RMSE compared to proposed G-Transformer in this paper, which uses the iterative training approach. One possible explanation is that during the simulation process, we need to dynamically model the outcomes and treatment variables. This simulation algorithm and our training algorithm share a similar step-by-step approach in term of generating results. Therefore, although both training methods attempt to optimize MSE loss over sequences of the same time step length, our iterative training approach can capture some biases introduced when using predicted outcomes and treatment variables as input for the prediction of next time step, thereby achieving better predictive performance.
Furthermore, as demonstrated in Table 5 and Figure 6, our experiments with the full Transformer architecture variant, the sequence-to-sequence G-Transformer featuring the full Encoder-Decoder architecture (Seq2seq G-Transformer + ED), did not perform as expected, possibly due to overfitting caused by an excessive number of parameters. Particularly over longer time intervals, the large number of parameters introduced by the decoder may have adversely affected the simulation process.
Table 5: CVSim: Counterfactual prediction under gc1. Performance of different variants of G-Transformer in predicting remaining 32-hour patient covariate trajectories based on first 34-hour patient covariate trajectories. Values reported represent the individual-level RMSE between the predicted and counterfactual trajectories. ED stands for using entire Transformer Encoder-Decoder rather than just using Transformer Encoder.
Individual Level
Model RMSEgc1
Seq2seq G-Transformer 1.550
Seq2seq G-Transformer + ED 1.935
Causal Transformer 1.435
Proposed G-Transformer 1.016
Figure 6: Individual-level RMSE over time under gc1 from different variants of G-Transformer
A.2 Gaussian Noise Simulation Experiment
In our proposed G-Transformer Monte Carlo simulation, we sample noise from a distribution of empirical loss from a held-out dataset. Here, we evaluate an alternative approach to sample noise from a normal distribution, with zero mean and empirical covariance estimated from data. As shown in Table 6, sampling noise empirically results in consistently better performance in term of individual-level RMSE across various models when applied to CVSim data under gc1.
Table 6: CVSim: Counterfactual prediction under gc1. Performance of various models with empirical and Gaussian noise sampling in predicting remaining 32-hour patient covariate trajectories based on first 34-hour patient covariate trajectories. Values reported represent the individual-level RMSE between the predicted and counterfactual trajectories.
Empirical Noise Gaussian Noise
Model Individual-level RMSEgc1 Individual-level RMSEgc1
Linear(g-comp) 1.103 1.418
G-Net 1.020 1.158
G-Transformer 1.016 1.240
A.3 Individual-Level RMSE Over Time
Table 7 presents the selected time step-specific individual-level RMSE across different models in CVSim experiment. G-Transformer generally performs better than linear implementation of g-computation and G-Net across the time steps.
Table 7: CVSim: Counterfactual prediction under gc1 and gc2. Performance of various models in predicting remaining 32-hour patient covariate trajectories based on first 34-hour patient covariate trajectories per time step. Values reported represent individual-level RMSE between the predicted and counterfactual trajectories.
Linear GNet GT
$t$ g-comp G-Transformer
gc1 1 0.688 0.654 0.615
8 1.062 1.020 1.023
16 1.125 1.049 1.042
24 1.168 1.052 1.047
32 1.255 1.106 1.092
gc2 1 0.791 0.752 0.720
8 1.122 1.079 1.077
16 1.208 1.131 1.123
24 1.286 1.188 1.165
32 1.348 1.213 1.186
Appendix B Additional MIMIC Experiments and Results
B.1 Adjustment of Individual-Level RMSE
Individual-level RMSE is adjusted to account for error due to over- and under-prediction in the trajectory length as a result of death and discharge. We use the same technique as in Su et al. to calculate the individual-level RMSE. Specifically, to account for potential over or under-predictions in trajectory length compared to ground truth, we adopt the following approach to modify the individual-level RMSE calculation. First, for patients who died within the first 24 hours, we replace all subsequent time steps of continuous variables with zero (normalized) until the 24th hour. Covariates with potential negative values are filled with the normalized minimum value in the dataset, while covariates following a log-normal distribution are filled with a normalized logarithmic zero. Second, for patients who were released from the hospital before the end of the first 24 hours in the ICUs, we fill all time steps following hospital release with the population-mean for all continuous variables (up until hour 24). Third, we do the same for the predicted chains (after the predicted death and release outcomes during the first 24 hours) so that all chains have lengths of 24-hours. We then compute the average RMSE across all patients by comparing the predicted trajectory (averaged across 100 Monte Carlo simulations per patient) and the ground-truth trajectory of each patient across all continuous variables.
Appendix C CVSim Data Generation
A CVSim 6-compartment circulatory model takes as inputs 28 variables that together govern a hemodynamic system. It then deterministically simulates forward in time a set of 25 output variables according to a collection of differential equations (parameterized by the input variables) modeling hemodynamics. Important variables in CVSim include arterial pressure (AP), central venous pressure (CVP), total blood volume (TBV), and total peripheral resistance (TPR). In real patients, physicians observe AP and CVP and seek to keep them above a clinically safe threshold. They do this by intervening on TBV (through fluid administration) and TPR (through vasopressors).
We defined simulated treatment interventions that were designed to mimic the impact of fluids and vasopressors. These simulated interventions alter the natural course of the simulation by increasing either TBV (in the case of the simulated fluids intervention) or TPR (in the case of the simulated vasopressor intervention). We generated patients by randomly initiating baseline inputs (which we hid from our G-Nets to make this a stochastic modeling problem) within plausible physiologic ranges, then using CVSim to simulate covariates forward, intervening according to the relevant treatment strategy at each timestep. Full details of the simulation process can be found in the Appendix.
Under (stochastic) observational treatment strategy $g_{o}$, the probability of receiving a non-zero vasopressor or fluid dose at a given time increases as MAP and CVP decrease according to a logistic regression function. Given that a dose is non-zero, the exact amount is drawn from a normal distribution with mean inversely proportional to MAP and CVP. Since all drivers of treatment under $g_{o}$ are observed in our data, the sequential exchangeability assumption holds and g-computation may be validly applied.
$g_{c1}$ is similar to $g_{o}$, except it is a deterministic treatment strategy and the functions linking treatment and dose to covariates have different coefficents. Under $g_{c2}$, treatment is always withheld.
Appendix D MIMIC Data and Experiments
Our cohort was limited to ICU stays in which the patient was identified as septic under the Third International Consensus Definitions for Sepsis and Septic Shock (Sepsis-3) Singer et al. (2016). Patients with missing records of pre-ICU fluids or admitted to the ICU following cardiac, vascular, or trauma surgery were removed.
We use AUC to evaluate our 24 hour outcome prediction performance. For each delay $k$, we do not include patients who have experienced death or release before hour $k$. A patient is positively labelled if they experience an outcome from hours $k$ to 24. Table 8 presents these test cohort statistics.
Table 8: 24-Hour Cohort Statistics. Size of the test sets used to evaluate AUCs and the respective percent of patients experiencing each outcome from hour $k$ to 24.
k Edema MV Diuretics Dialysis Release Death
2 (894, 29.98%) (894, 40.94%) (894, 15.10%) (894, 2.68%) (894, 13.20%) (894, 2.35%)
4 (894, 29.87%) (894, 40.04%) (894, 13.87%) (894, 2.68%) (894, 13.20%) (894, 2.35%)
6 (894, 29.75%) (894, 38.81%) (894, 12.75%) (894, 2.68%) (894, 13.20%) (894, 2.35%)
8 (892, 29.82%) (892, 36.77%) (892, 12.22%) (892, 2.69%) (892, 13.12%) (892, 2.24%)
As predictors to our model, we selected covariates that are typically monitored in the ICU and important for determining sepsis intervention strategies, as well as potential confounders. The covariates we used were similar to that of Li et al. Li et al. (2021), encompassing, but not limited to, basic demographic information, an Elixhauser comorbidity score, a SOFA score, laboratory values and vital signs, and urine output Li et al. (2021). A comprehensive list is provided in Tables 9 and 10. The demographics, comorbidities, and pre-ICU fluids were unmodeled and regarded as static while the remaining variables were modeled and regarded as dynamic (time-varying).
Table 9: MIMIC static variables. All variables were used as inputs to our models.
Variable Name Variable Type Units
Age Continuous years
Gender Binary N/A
Pre-ICU Fluid Amount Continuous mL
Elixhauser Score Continuous N/A
End Stage Renal Failure Binary N/A
Congestive Heart Failure Binary N/A
Table 10: MIMIC time-varying variables. All variables were used as inputs to our models, and boluses and vasopressors were also intervention variables. *Refers to maintenance fluids (not an intervention).
Variable Name Variable Type Units
Heart Rate Continuous beats/min
Diastolic Blood Pressure Continuous mmHg
Systolic Blood Pressure Continuous mmHg
Mean Blood Pressure Continuous mmHg
Minimum Mean Blood Pressure Continuous mmHg
Minimum Change in Mean Blood Pressure from Baseline Continuous mmHg
Minimum Mean Blood Pressure from Baseline Continuous mmHg
Minimum Change in Mean Blood Pressure from Previous Continuous mmHg
Temperature Continuous degree C
SOFA Score Treated as Continuous N/A
Change in SOFA Score from Baseline Treated as Continuous N/A
Change in SOFA Score from Previous Treated as Continuous N/A
Platelet Continuous counts/$10^{9}$
Hemoglobin Continuous g/dL
Calcium Continuous mg/dL
BUN Continuous mmol/L
Creatinine Continuous mg/dL
Bicarbonate Continuous mmol/L
Lactate Continuous mmol/L
O2 Requirement Level Continuous N/A
Change in O2 from Baseline Continuous N/A
Change in O2 from Previous Continuous N/A
pO2 Continuous mmHg
sO2 Continuous %
spO2 Continuous %
pCO2 Continuous mmHg
Total CO2 Continuous mEq/L
pH Continuous Numerical[1,14]
Base excess Continuous mmol/L
Weight Continuous kgs
Change in Weight Continuous kgs
Respiratory Rate Continuous breaths/min
Fluid Volume* Continuous mL
Urine Output Continuous mL
Cumulative Edema Binary N/A
Pulmonary Edema Indicator Binary N/A
Diuretics Indicator Binary N/A
Dialysis Indicator Binary N/A
Mechanical Ventilation Indicator Binary N/A
Vasopressor Indicator Binary N/A
Bolus Volume Continuous mL
In-Hospital Mortality Indicator Binary N/A
Release Indicator Binary N/A
Cumulative Fluids Continuous mL
Fluid Balance in ICU Continuous mL
Appendix E Hyperparameter Settings
E.1 Settings in Cancer Growth Experiments
Hyperparameter optimization on Cancer Growth data was performed by searching over the hyperparameter space shown in Table 11 with optimal parameters starred. The experiments were performed on NVIDIA Tesla K80 GPU with 12 vCPUs + 110 GB memory. Hyperparameter settings of rMSN and CRN were as reported in Li et al. (2021).
Table 11: Hyperparameter search space in Cancer Growth experiments.
Hyperparameters Search Range
Linear (g-comp) Learning Rate 0.001, 0.01
Number of Layers (Representation) 1, 2
Hidden Dimension (Representation) 8, 16, 32, 64
G-Net Learning Rate 0.001, 0.01
Number of Layers (Continuous) 1, 2, 4
Hidden Dimension (Continuous) 4, 16, 64
Number of Layers 2, 4, 6
Hidden Dimension 32, 64, 128
Batch Size 16, 32
G-Transformer Learning Rate 0.0001
E.2 Settings in CVSim Data
[…]
Table 12: Hyperparameter search space in CVSim experiments.
Hyperparameters Search Range
Batch Size 16
Linear (g-comp) Learning Rate 0.0001
Number of Layers (Continuous) 2, 3
Hidden Dimension (Categorical) 64, 128
Hidden Dimension (Continuous) 64, 128
Batch Size 16
G-Net Learning Rate 0.0001
Number of Layers 3
Hidden Dimension 32, 64, 128
Batch Size 16
G-Transformer Learning Rate 0.0001
E.3 Settings in MIMIC Data
[…]
Table 13: Hyperparameter search space in MIMIC experiments.
Hyperparameters Search Range
Number of Layers (Continuous) 2, 3
Hidden Dimension (Categorical) 64
Hidden Dimension (Continuous) 64, 128
Batch Size 16
Weight Decay 1e-6
G-Net Learning Rate 0.0001
Number of Layers 2, 3
Hidden Dimension 64, 128
Batch Size 16
Weight Decay 1e-5
G-Transformer Learning Rate 0.0001
E.4 Settings for Causal Transformer
Different from the Causal Transformer paper, we did not generate different counterfactual trajectories on the Cancer Growth dataset at each time step using Single sliding treatment or Random trajectories. For each patient in the test set, we inserted our counterfactual treatment at the last 5 time steps and applied this counterfactual treatment in each subsequent time step. In addition, for multi-step time prediction with Causal Transformer, we predicted the next 4 time steps and $\tau=5$ in this dataset. In both the CVSim dataset and the MIMIC dataset, since the model setup does not allow access to the actual treatment, we modified parts of the Causal Transformer’s code to enable the model to use the predicted treatment as input for predicting the outcomes of the next time step.
Hyperparameter Settings For MIMIC and Cancer growth dataset, given that the Causal Transformer has already provided rich information on hyperparameter settings in the original paper (Melnychuk et al., 2022), we used their hyperparameter settings in experiments. For CVSim dataset, we made slight adjustments to the hidden dimension for experimentation as its format is similar to that of the MIMIC dataset. The specific hyperparameter settings are shown in the table 14.
Table 14: Hyperparameter setting for Causal Transformer in different experiments.
Hyperparameters Search Range
Dropout rates 0.2
Number of Layers 2
Number of attention heads 3
Number of epochs 100
Hidden Dimension 44
MIMIC Learning Rate 0.001
Dropout rates 0.1
Number of Layers 2
Number of attention heads 3
Number of epochs 150
Hidden Dimension 48
Cancer Growth Learning Rate 0.0001
Dropout rates 0.2
Number of Layers 2
Number of attention heads 3
Number of epochs 100
Hidden Dimension 24
CVSim Learning Rate 0.0001